What mass of Oxygen gas would occupy a volume of 1800mL at Standard conditions? Atomic weight of oxygen is 16.
1
Expert's answer
2012-06-05T11:26:08-0400
Standart conditions:

Temperature 273.15 K

Pressure 101.325 kPa

Using the universal gas law:

pV = nRT

or

pV = mRT/M

hence,

m(O2) = pVM/RT

where

p - pressure

V - volume

M - molar mass of O2

R - the universal gas law constant of 8.3145 m^3*Pa/(mol*K)

M(O2) = 2*16 = 32 g/mol

m(O2) = 101325 Pa * 0.0018 m^3 * 32 g/mol / ( 8.3145 m^3*Pa/(mol*K) * 273.15 K) = 2.57 g